OpenStudy (anonymous):

Solve the following system of equations: 3x - 2y = 6 6x - 4y = 14 (0, 0) (6, 14) Infinitely many solutions No solutions

OpenStudy (anonymous):

Ill medal and fan anybody who answers! :)

hero (hero):

If you're only interested in answers, you should try Yahoo Answers. Here, we "study", explain, guide, and help users understand the process of how to solve the problem on their own.

OpenStudy (anonymous):

I meant answer it with me.

hero (hero):

Hint: Divide both sides of the second equation by 2

OpenStudy (anonymous):

Ok so 3x-y=3

hero (hero):

Double check your work.

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

@VortexAlliby

hero (hero):

If you divided both sides of the second equation by two correctly, you would end up with the system: 3x - 2y = 6 3x - 2y = 7

OpenStudy (anonymous):

Ok. I get that now. lol I am a complete moron.